OpenStudy (anonymous):

prove for each integer n, n is even if and only if 4 divides n^2

OpenStudy (anonymous):

hint a general even number looks like \(n=2j\) for some \(j\)

OpenStudy (anonymous):

I am having trouble....

OpenStudy (anonymous):

better yet what I am having trouble is proving the contrapositive of: if n is odd then 4 does not divide n^2.

OpenStudy (anonymous):

Just like Satellite wrote the general form of an even number, what is the general form of an odd number? What happens when you square it?

OpenStudy (anonymous):

I got the first part which is: for each integer n, if n is even, then 4 divides n^2.

OpenStudy (anonymous):

nerdguy2535, the def of an odd number is 2k+1, I already worked it out which (2k+1)^2=4k^2+4k+1, then I factored it

OpenStudy (anonymous):

and got 2(2k^2+2k)+1, therefore n^2=2q+1 for some q is element of integers

OpenStudy (anonymous):

Nice. Is it possible for an odd number to be divisible by 4?

OpenStudy (anonymous):

no, I guess what I am not seeing is how to make the connection in the end in words, to finalize my proof...

OpenStudy (anonymous):

why go that route? you can do it directly much faster and mostly direct proofs are nicer and more convincing

OpenStudy (anonymous):

takes only one or two lines in each direction

OpenStudy (anonymous):

He's already done doing the contrapositive though.

OpenStudy (anonymous):

satellite73, I tried proving it directly but it got messy with the square root, because original section of proof is: if 4 divides n^2, then n is even. Unless I was doing it wrong?

OpenStudy (anonymous):

for example if \(n\) is even, then \(n=2j\) for some integer \(j\) making \[n^2=(2j)^2=4j^2\] divisible by \(\) done

OpenStudy (anonymous):

ooh i see

OpenStudy (anonymous):

there is always the fundamental theorem of arithmetic but i guess the contrapositive works nicely without it

OpenStudy (anonymous):

what is the fundamental theorem of arithmetic?